mr. avila is for times as old as his son. in 16 years he will be only twice as old. what is the present age of each?

Answer by NeedHelpUrgently(2)

Dad = 4x & Son = x Now.
In 16 years:
4x+16=2(X+16)
4x+16=2X+32
4x-2X=32-16
2X=16
x=8
X=8 Age of son now.
4*8=32 Age of dad now.
Proof:
4*8+16=2(8+16)
32+16=2*24
48=48